KyledaGreat:

Determine the more basic function that has been shifted, reflected, stretched, or compressed.

KyledaGreat:

\[v(x) = -2\sqrt{x-1} + 3\] \[f(x) =\]

KyledaGreat:

\[v(x) = -2\sqrt{x-1} + 3\] \[f(x) =\]

KyledaGreat:

@extrinix

KyledaGreat:

@minustempo

minustempo:

The basic function is what is left when you talk away all constants and coefficients.

minustempo:

*take away

KyledaGreat:

The basic function is x^3 (-2-x)^3 = -(x-1) + 3 is this right ?

KyledaGreat:

@extrinix

KyledaGreat:

@vocaloid @tranquility @surjithayer

surjithayer:

\[f(x)=\sqrt{x}\]

KyledaGreat:

1 attachment
KyledaGreat:

1 attachment
KyledaGreat:

1 attachment
KyledaGreat:

1 attachment
surjithayer:

e